OpenStudy (whalenjacob):

Part A) 5x+3(x−5)=6x+2x−15 Simplify both sides of the equation. 5x+3(x−5)=6x+2x−15 Simplify: 5x+3(x−5)=6x+2x−15 5x+(3)(x)+(3)(−5)=6x+2x+−15(Distribute) 5x+3x+−15=6x+2x+−15 (5x+3x)+(−15)=(6x+2x)+(−15)(Combine Like Terms) 8x+−15=8x+−15 8x−15=8x−15 Subtract 8x from both sides. 8x−15−8x=8x−15−8x −15=−15 Add 15 to both sides. −15+15=−15+15 0=0 Only one solution. Part B) I used distributive property.

OpenStudy (whalenjacob):

anyone? i need help

OpenStudy (whalenjacob):

@mathmale @mathmate

OpenStudy (whalenjacob):

can you tell the answer then explain??

OpenStudy (evoker):

That first part of 0=0 means there are an infinite number of solutions.

OpenStudy (evoker):

one solution would be something like x=4, while no solutions would be something like 0=10

OpenStudy (evoker):

As for part B what is the question?

OpenStudy (whalenjacob):

How many packets of chocolate wafers and vanilla wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer.

OpenStudy (whalenjacob):

there you go

OpenStudy (evoker):

Um what was the initial problem I guess then, all I see is the algebraic equation.

OpenStudy (whalenjacob):

here ill give you the whole question

OpenStudy (whalenjacob):

Josh and his friends bought vanilla wafers for $4 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $45 to buy a total of 27 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of vanilla wafers and the number of packets of chocolate wafers that Josh and his friends bought at the carnival. Define the variables used in the equations. Part B: How many packets of chocolate wafers and vanilla wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer.

OpenStudy (evoker):

Ah ok so you would have two equation lets call V the number of vanilla wafers and C the number of chocolate wafer

OpenStudy (whalenjacob):

whalenjacob is typing...

OpenStudy (whalenjacob):

ok

OpenStudy (evoker):

One equation for the total spent, and another for the total number of items.

OpenStudy (whalenjacob):

ok

OpenStudy (evoker):

So for the number of items V+C=27

OpenStudy (evoker):

Since they bought a total of 27 items

OpenStudy (whalenjacob):

alright

OpenStudy (evoker):

For money spent 4V+1C=45

OpenStudy (whalenjacob):

ok

OpenStudy (evoker):

So now we use either substitution or elimination method for two equations.

OpenStudy (evoker):

Do you have a preference

OpenStudy (whalenjacob):

not really anyone you choose im fine with

OpenStudy (evoker):

Ok rearranging the first equation we get V=27-C

OpenStudy (whalenjacob):

yup[

OpenStudy (evoker):

Then plugging into the second we have 4(27-C)+1C=45

OpenStudy (evoker):

So now we can solve for C

OpenStudy (whalenjacob):

ok

OpenStudy (evoker):

So using the distributive property that become 4*27-4C+1C=45

OpenStudy (whalenjacob):

ok i got it

OpenStudy (evoker):

Ok so what do you get for C

OpenStudy (whalenjacob):

21?

OpenStudy (evoker):

Thats what I get

OpenStudy (evoker):

So finally since you know there are a total of 27 items the remaining are vanilla.

OpenStudy (whalenjacob):

i thought it was 21

OpenStudy (evoker):

Right 21 Chocolate and the remainder of the 27 candies are vanilla.

OpenStudy (whalenjacob):

oh duh im sorry

OpenStudy (evoker):

So 21 Chocolate and 6 Vanilla

OpenStudy (evoker):

And that would be one way of solving the problem.

OpenStudy (whalenjacob):

ok ill be right back stay and ima be right back